If the demand law is p = 85 - 4x– x2. what will be the consumer's surplus if : (i) xo =5 and (ii) po = 64
1
Expert's answer
2019-09-30T09:11:23-0400

The surplus is an area below demand function and above the line p=64. So we need to calculate the integral on [0; 5]:

"\int85-4x-x^2-64 = 21x - 2x^2 - \frac{x^3}{3} =
21*5-2*5^2 - \frac{5^3}{3} = 13.33(33)"